OpenStudy (anonymous):

Marty's class collected $32.00 in a school-wide penny collection contest. If Tommy's class added $6.00 to the amount they collected, x, and then multiplied the sum by 2, they would have the same amount of money as Marty's class. 2(x + $6.00)= $32.00 How much money did Tommy's class collect? @apoorvk

OpenStudy (alexandervonhumboldt2):

2x+12=32 solve this

OpenStudy (alexandervonhumboldt2):

expnad ( ) using disriburive property a*(b+c)=ab+ac